OpenStudy (anonymous):

(Sin2θ+sin4θ/sin2θ-sin4θ)+(Tan3θ/tanθ)=0

OpenStudy (loser66):

\[\dfrac{sin(2\theta)+ sin(4\theta)}{sin(2\theta)-sin(4\theta)}+\dfrac{tan(3\theta)}{tan(\theta)}=0\] is it right?